Find values for m and b in the following system so that the solution to the system is (-3, 4).
5x + 7y = b
mx + y = 22
Please help me with this.
Thank you Debbie

Answer by edjones(8007)

Find values for m and b in the following system so that the solution to the system is (-3, 4).
I assume x=-3 and y=4
5x + 7y = b
mx + y = 22
.
5*-3+7*4=b
-15+28=b
b=13
.
-3m+4=22
-3m=18 subtract 4 from each side.
m=-6 divide each side by -3.
Ed